OpenStudy (anonymous):

I have 9 math questions. Can i have some help?

OpenStudy (anonymous):

what type of math is it?

OpenStudy (anonymous):

Algebra

OpenStudy (anonymous):

whats the questions

OpenStudy (anonymous):

okay yea im pretty good at algebra

OpenStudy (anonymous):

I can try what grade are you in?

OpenStudy (anonymous):

Jill has been watching the squirrels that live in her neighborhood. The number of squirrels change each week  Which function best shows the relationship between n and f(n) A. f(n) =12(1.5)n B. f(n) = 12(1.5)n-1 C.f(n) = 8(1.5)n D. f(n) = 8(1.5)n-1

OpenStudy (anonymous):

@izzy529 im in 9th grade

OpenStudy (anonymous):

@damneny help? ^

OpenStudy (anonymous):

Oh sorry I'm in 7th grade :(

OpenStudy (anonymous):

that is fine @izzy529

OpenStudy (anonymous):

@damneny can i have your help? You really helped me last time and i could use your brain again

OpenStudy (anonymous):

The correct answer would be D. f(n) = f(1) • r^n-1 8*1.5^4-1 8*1.5^3 = 27.

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

Anything else?

OpenStudy (anonymous):

Not now but when i do i will let you know. :)

OpenStudy (anonymous):

Sure :>

OpenStudy (anonymous):

E.J has shown that a function, f(x) , grows by 12% for every unit in the domain. What does this prove? A. The function f(x) is an arithmetic sequence B. The function f(x) is a geometric sequence C. The function f(x) is not a sequence D. This does not prove anything

OpenStudy (anonymous):

@theopenstudyowl help?

OpenStudy (anonymous):

@confluxepic help? ^^